4. The daily amount of red bull, in liters, dispensed by a machine is a random variable X having a continuous uniform distribution with 4< x< 8.

Find the probability that on a given day the amount of red bull dispensed by this machine will be.

More than 7 liters or less than 5 liters More than 6.4 liters

What is the average daily amount of red bull in liters dispensed by this machine
Answer #1

X ~ U(4, 8)

A) P(X > 7 or X < 5) = P(X > 7) + P(X < 5)

= (8-7)/(8-4) + (5-4)/(8-4)

= 0.5

B) P(X > 6.4) = (8-6.4)/(8-4) = 0.4

C) average = (4 + 8) / 2 = 6